OpenStudy (anonymous):

Check work? st + 3t = 6 solve for s

OpenStudy (anonymous):

st + 3t = 6 divide by s on both sides t + 3t = 6/s multiply both sides by 6 6t + 18t = s ?

OpenStudy (anonymous):

or would it be: 6t + 32 = s?

OpenStudy (anonymous):

st + 3t = 6 Subtract 3t st = 6 - 3t Divide by t s = (6 - 3t)/ t

OpenStudy (anonymous):

thanks you guys!